Circular No. 6893

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998bq, 1998br, 1998bs L. Germany reports that the Mount Stromlo Abell Cluster Supernova

Search Team (cf. IAUC 6639) has discovered three apparent supernovae on V and R CCD images taken with the Mount Stromlo 1.27-m telescope (+ Macho Camera):

SN  1998 UT  α_{2000}  δ_{2000}  V  Offset
1998bq Apr. 28 15^{h}10^{m}20.^{s}53 + 5^{o}51′16′′.5 19.2 1′′.2 W, 1′′.7 S
1998br Apr. 18 15 10 24.94 + 5 55 30.0 20.5 5′′.1 E, 3′′.6 N
1998bs Apr. 18 13 26 31.51 −31 09 08.9 20.6 0′′.1 E, 0′′.5 N

SNe 1998bq and 1998br are near the cluster Abell 2029; SNe 1998bs is near Abell 3558. These object were not visible on images obtained on the following dates (with limiting V magnitudes given): SN 1998bq, Apr. 18, 21.5; SN 1998br and SN 1998bs, Mar. 10, 21.0. The limiting magnitude on Apr. 28 was V = 20.5. Confirmation images for the other supernovae were taken with the same telescope on Apr. 30 (limiting mag V = 21.5). Finding charts

are available at http://msowww.anu.edu.au/~reiss/Abell SNSearch/.

KLEINMANN-LOW NEBULA T. Omodaka and T. Maeda, Department of Physics, Kagoshima Uni-

versity; and N. Mochizuki, National Astronomical Observatory of Japan, report: “A burst of water masers (rest frequency 22.235 GHz) in the Orion-KL region (α = 5^{h}32^{m}47^{s}.01, δ = −5^{o}24′23′′.8, equinox 1950.0) was detected in January with the Kagoshima 6-m radio telescope (+ radio Acousto Optical Spectrometer). The central velocity of the burst peak is 7.8 km/s. The intensity of the flare component has increased gradually since January and is now 138 times that of the ‘quiescent’ flux density of the water masers in this region, as indicated by the following peak flux densities (V_{LSR} = 7.8 km/s): 1997 Jan. 20, 13 000 Jy; Dec. 4, 18 000; 1998 Jan. 2, 40 000; Feb. 4, 72 000; Mar. 4, 200 000; Mar. 26, 870 000; Apr. 28, 1 800 000. We also confirmed the burst with the Nobeyama 45-m, Kashima 34-m, and Mizusawa 10-m telescopes. Because the beam size of Nobeyama 45-m telesope is 74′′, we are sure the burst is at the Kleinmann-Low nebula. An accurate position of the burst component will be obtained by VLBI observations.”

1998 May 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green